# Prioritized-MVBA: A New Approach to Design an Optimal Asynchronous Byzantine Agreement Protocol

### v1

@@ 4 The Proposed Protocol
In this section, we present a protocol for multi-valued Byzantine agreement capable of tolerating up to $f<\frac{n}{3}$ Byzantine faults. The protocol features an expected communication bit complexity of $O(ln^{2}+\lambda n^{2})$ and an expected constant asynchronous round count. Our modular implementation comprises five distinct sub-protocols: committee selection, prioritized verifiable consistent broadcast protocol (pVCBC), Propose-Recommend, random order/permutation generation, and Sequential-ABBA. The framework of the proposed protocol is illustrated in Figure 1(a). The violet-colored boxes indicate our contributions to the protocol. We break down the protocol into two main parts. The top part includes the committee selection and the VCBC protocol for the committee members, while the bottom part handles the distribution of the committee members’ messages and the agreement protocol. Section 4.1 provides an overview and the construction of party selection and broadcast (top part of the framework), and Section 4.2 offers a detailed exposition of the protocol, including pseudocode and descriptions.
4.1 Optimizing Message and Communication Overhead
To enhance the efficiency of Byzantine agreement protocols, we propose a novel method to minimize the number of broadcast requests. Rather than requiring all $n$ parties to broadcast their requests, we select a subset of parties to do so, ensuring that the protocol’s progress and security properties are maintained.
Party Selection and Broadcast: In traditional asynchronous Byzantine agreement protocols, all $n$ parties are typically allowed to broadcast their proposals. However, the main factor contributing to the $O(n^{3})$ communication complexity in classic MVBA protocols is that each of the $n$ parties must broadcast an $O(n)$-size array. These arrays include broadcast information from all parties. Our approach aims to reduce the number of broadcasting parties, thereby eliminating the need for transmitting an $O(n)$-size array and decreasing the number of required Asynchronous Binary Byzantine Agreement (ABBA) instances. From this point forward, we use proposal instead of request and propose instead of broadcasting requests.
To achieve this reduction, we select at least $f+1$ parties to broadcast their proposals, ensuring that there is at least one honest party among them. This selection is crucial for guaranteeing the protocol’s progress. We employ a committee selection protocol similar to FasterDumbo [11], which uses a cryptographic coin-tossing scheme to randomly select $\kappa$ (a security parameter) parties. We set $\kappa=f+1$ to ensure the inclusion of at least one honest party, increasing the likelihood of multiple honest parties being selected.
Dynamic selection of parties for each protocol instance prevents adversaries from corrupting fixed parties, ensures fair opportunity for all parties to propose, and mitigates the risk of starvation and Denial-of-Service (DoS) attacks. The committee selection process must be immune to adversary influence. The cryptographic coin-tossing scheme ensures random selection, requiring honest party participation after the completion of the previous instance. We used the committee member/selected party alternatively.
(a) Committee-Selection protocol. Each party shares their $s_{i}=CShare(r_{i})$ and a party requires f+1 shares. We see a party use $CToss(r_{1},s)$ function to get the selected parties and each party gets the same set of parties. We show only for the two parties but it is true for every party.
(b) pVCBC protocol. Each selected party (here $p_{1}$ and $p_{2}$ ) propose its requests. Each receiving party verify the requests and add its sign-share on the message and returns to the sender. A party waits for 2f+1 sign-shares (here it is 3 because f = 1). The 2f+1 sign shares are necessary to create a threshold-signature ($ts1$).
Committee Selection Protocol: The committee selection protocol (CS) is an integral part of our approach. It ensures that a reliable subset of parties is chosen to broadcast proposals, thus reducing communication complexity. The steps of the CS protocol are illustrated in Algorithm 1 and described as follows:
@@ 4.1 Optimizing Message and Communication Overhead
  • Generating Coin-Shares: Upon invocation, a party generates a coin-share $\sigma_{i}$ for the current instance and broadcasts it to all parties, waiting to receive $f+1$ coin-shares. (lines 3-5)
@@
  • Verifying Coin-Shares: When a party receives a coin-share for the first time from a party $p_{k}$, it verifies the coin-share and accumulates it in set $\Sigma$. This process continues until $f+1$ valid shares are collected. (lines 8-10)
@@
  • Selecting Parties: After receiving $f+1$ valid coin-shares, a party uses the $CToss$ function along with the collected coin-shares to select $f+1$ parties. (lines 6-7)
@@ 3      $\Sigma\leftarrow\{\}$
4
5
6      upon $CS\langle id,instance\rangle$ invocation do
7           $\sigma_{i}$leftarrow $CShare\langle r_{i}\rangle$
8           $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
9           wait until $|\Sigma|=f+1$
10           return $CToss\langle r_{i},\Sigma\rangle$
11
12
@@ 8           $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
15                     $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
16
@@ 15                     $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
Algorithm 1 Committee Selection Protocol for party $p_{i}$
We provide a illustration of the CS protocol in figure 2(a) and the cryptographic coin-tossing definitions (e.g. $CShareVerify$) in the Appendix A.3.
Broadcast protocol A committee member requires a verifiable proof of its proposal to demonstrate that it has broadcast the same proposal to at least $f+1$ honest parties. Typically, the VCBC is employed for this purpose, allowing every party to generate a verifiable proof. Since our protocol restricts proposal broadcasting to only the selected parties, we employ a slightly modified version of the verifiable consistent broadcast protocol, termed pVCBC. This adaptation ensures that if a party receives a verifiable proof from another party, it does not need to verify whether the proof originates from a selected party or not. The definition of the VCBC protocol and the construction of the pVCBC protocol can be found in the Appendix A.1 and Appendix C. We have also provided an illustration of the protocol in Figure 2(b).
4.2 Prioritized MVBA
This section details the MVBA module in an asynchronous network. We propose a novel MVBA protocol called Prioritized-MVBA (pMVBA), which uses a committee as a prioritized parties and reduces the communication complexity of the classic MVBA protocol.
Figure 3: New-VBAConst protocol. After completing the pVCBC broadcast, each party proposes its value along with verifiable proof. Upon receiving verifiable proof, the party recommends the received proposal. Each party waits for $n-f$ recommendations, then uses the coin-tossing scheme (CShare) to generate a permutation and initiates the ABBA protocol to reach an agreement on one of the proposals.
Overview of pMVBA The pMVBA protocol is inspired by the asynchronous binary Byzantine agreement methodology from Cachin et al. [13]. It significantly reduces the number of initial broadcasts, eliminating costly sub-components by leveraging the message distribution pattern of the adversary.
In classic MVBA, each party broadcasts an $O(n)$-size array to achieve $O(1)$ expected constant runtime. Our pMVBA protocol as shown in Fig 3 and explained in Section 4.1, however, restricts this broadcasting to $f+1$ parties, significantly reducing communication complexity. The protocol replaces the $O(n)$ commit broadcast with a single recommend message. We provide a details about classic MVBA protocol challenges in the Appendix B.2.
Since $f+1$ parties generate verifiable proofs and broadcast their proposals, a party must respond after receiving the first verifiable proof. If multiple parties complete the pVCBC protocol, different parties may receive proofs from different parties, hindering agreement. Therefore, we introduce the recommend step, where a party broadcasts received proofs to all parties, ensuring at least one proof reaches the majority of the parties ($12f+$ 1), which is necessary to employ ABBA protocol to reach an agreement. We use majority and threshold-number alternatively.
Replacing Commit with Recommend The key idea in pMVBA is to replace the commit step with the recommend step. The commit step in classic MVBA protocols involves collecting $n-f$ verifiable proofs and creating an array of length $n$. In contrast, the recommend step involves broadcasting verifiable proofs directly, significantly reducing complexity.
Construction of the Recommend Step The recommend step consists of sending and collecting recommendations. This step ensures that at least one verifiable proof reaches the majority of parties. The detailed SCR protocol is shown in Algorithm 2:
  • Initialization: A party initializes an empty set $\Sigma$.
  • Broadcast Recommendation: A party creates a recommendation message and broadcasts it to all parties.
  • Recommendation Collection: Upon receiving a recommendation, a party verifies the threshold-sign and instance. If valid, the party adds the proposal and $\sigma$ to its set $\Sigma$.
  • A party waits for $n-f$ recommendation messages before proceeding.
1 Local variables initialization:
2 $\Sigma\leftarrow\{\}$
3
4 upon $SCR\langle ID,proposal,\sigma\rangle$ invocation do
[…]
6       wait until $|\Sigma|$ = $n-f$
7       return $\Sigma$
8
9
10

### v2

@@ 4 Design of pMVBA
In this section, we present a protocol for multi-valued Byzantine agreement capable of tolerating up to $f<\frac{n}{3}$ Byzantine faults. The protocol features an expected communication bit complexity of $O(ln^{2}+\lambda n^{2})$ and an expected constant asynchronous round count. Our modular implementation comprises five distinct sub-protocols: Committee Selection (CS), Prioritized Verifiable Consistent Broadcast (pVCBC), Propose-Recommend, Random Order/Permutation Generation, and Sequential-ABBA. The framework of the proposed protocol is illustrated in Figure LABEL:fig:Framework.
4.2 Committee Selection
In the proposed pMVBA protocol, the CS sub-protocol plays a critical role in optimizing efficiency by reducing the communication complexity inherent in Byzantine agreement protocols. Rather than requiring all $n$ parties to broadcast their proposals, our method strategically selects a subset of at least $f+1$ parties, based on the party ID and the current instance of the protocol, to perform this task. This selection ensures the protocol’s progress while maintaining its security properties, with the output being a set of $f+1$ parties designated as committee members for that specific instance.
To achieve this, the CS protocol utilizes a cryptographic coin-tossing scheme, inspired by Dumbo [FASTERDUMBO], to dynamically and randomly select $\kappa=f+1$ parties for each protocol instance. This selection process guarantees that at least one honest party is included in the committee, thereby preserving the integrity of the protocol.
The dynamic nature of party selection mitigates several risks, including adversarial corruption, ensuring fair participation across all parties and reducing the likelihood of starvation and Denial-of-Service (DoS) attacks. The randomness introduced by the cryptographic coin-tossing scheme is crucial in preventing adversary influence and securing the CS process.
The CS protocol is illustrated in Algorithm 1 and involves the following steps:
@@ 4.2 Committee Selection
  • Generating Coin-Shares: Upon invocation, each party generates a coin-share $\sigma_{i}$ and broadcasts it to all parties. The party then waits to receive $f+1$ coin-shares. (lines 3-5)
@@
  • Verifying Coin-Shares: Upon receiving a coin-share from another party $p_{k}$, the receiving party verifies the coin-share and adds it to the set $\Sigma$ until $f+1$ valid shares are collected. (lines 8-10)
@@
  • Selecting Parties: After collecting $f+1$ valid coin-shares, the $CToss$ function is used to determine the selected subset of parties. (lines 6-7)
@@
2Local variables initialization:
3   $\Sigma\leftarrow\{\}$
4   $CUR\_INSTANCE\leftarrow 0$
5
6
7   upon $CS\langle ID\rangle$ invocation do
8     $CUR\_INSTANCE\leftarrow ID.INSTANCE$
9     $id\leftarrow ID.id$
10     $\sigma_{id}$ $\leftarrow$ $CShare\langle id\rangle$
11
12    send $\langle SHARE,id,\sigma_{id},CUR\_lINSTANCE\rangle$ to all parties
13     wait until $|\Sigma|=f+1$
14     return $CToss\langle id,\Sigma\rangle$
15
16
18       if CUR_INSTANCE=INSTANCECUR\_INSTANCE=INSTANCE and $CShareVerify\langle id,k,\sigma_{k}\rangle=true$ then
19         $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
20
21
Algorithm 1 Committee Selection for party $p_{i}$
(a) Committee-Selection protocol. Each party shares their $s_{i}=CShare(r_{i})$, and a party requires $f+1$ shares. The $CToss(r_{1},s)$ function is used to select the parties, ensuring that each party gets the same set. This illustration shows the process for two parties, but it applies universally.
(b) pVCBC protocol. Each selected party (here $p_{1}$ and $p_{2}$ ) proposes its requests. Each receiving party verifies the requests, adds its sign-share to the message, and returns it to the sender. A party waits for 2f+1 sign-shares (here it is 3 because f = 1). The 2f+1 sign shares are necessary to create a threshold-signature ($ts1$).
Figure 3: Committee Selection and pVCBC protocol
The CS protocol, as depicted in Figure 3(a), ensures a secure and efficient selection of broadcasting parties, which significantly reduces the overall communication complexity of the pMVBA protocol. Definitions of the cryptographic coin-tossing functions (e.g., CShareVerifyCShareVerify) are provided in Appendix A.3.
4.3 pVCBC
Following the selection of committee members through the Committee Selection protocol, each selected member must provide a verifiable proof of their proposal to ensure that it has been broadcast to at least $f+1$ honest parties. Specifically, the input for this protocol includes the ID, requests, and the selected parties, while the output is a threshold-signature—a verifiable proof that the same requests have been sent to at least $f+1$ honest parties. This proof is crucial for maintaining the integrity and consistency of the protocol, as it verifies that the proposal has been correctly disseminated among the parties. Traditionally, the Verifiable Consistent Broadcast (VCBC) protocol is used to generate such proofs, enabling every party to produce a verifiable record of their broadcast proposals. The definition of verifiability and the VCBC protocol itself are detailed in Appendix A.1. However, since our protocol restricts broadcasting to only the selected committee members, we utilize a slightly modified version of the VCBC protocol, which we term pVCBC (Prioritized Verifiable Consistent Broadcast).
The pVCBC protocol aligns with the selective broadcasting approach from the Committee Selection process. This adaptation ensures that when a party receives verifiable proof from another, there’s no need to verify its origin from a selected committee member, as the protocol inherently guarantees this. This streamlines verification and reduces unnecessary checks, preserving the efficiency introduced by Committee Selection. The pVCBC protocol’s construction is detailed in Algorithm 2, and an illustration is provided in Figure 3(b), visually representing its interactions and verification steps as a critical component of the overall pMVBA protocol.
Construction of the pVCBC
  • Upon invocation of the pVCBC protocol, a party sends (ID, requests, SelectedParties) to the parties. (lines 03).
  • Upon receiving a message (ID, requests, SelectedParties) from a party $p_{k}$, a party checks whether the sender is a selected party. If the sender is a selected party, the party adds its signature share (signsharesign-share) to the message, resulting in $\sigma$. The party then replies with $\sigma$ to the sender (lines 11-14).
  • Upon receiving a signature share $\sigma_{k}$ from a party $p_{k}$, a selected party adds the signature share $\sigma_{k}$ to its set $\Sigma$ (lines 07 - 09).
  • A selected party waits for $\langle n-f\rangle$ valid signature shares. Upon receiving nfn-f signature shares, the party combines them to generate a threshold signature $\rho$ (proof that the party has sent the same request to at least $f+1$ honest parties) and returns $\rho$ to the caller (line 05).
1
[…]
3   $\Sigma\leftarrow\{\}$
4   send $\langle ID,requests\rangle$ to all parties
5   wait until $|\Sigma|=n-f$
6   return $\rho\leftarrow CombineShare_{id}\langle requests,\Sigma\rangle$
7
8
9
[…]
11     if $VerifyShare_{k}\langle requests,(k,\sigma_{k})\rangle$ then
12       $\Sigma\leftarrow{\Sigma\cup\{\sigma_{k}\}}$
13
14
15       upon receiving $\langle ID,requests\rangle$ from the party $p_{j}$ for the first time do
16         if $p_{j}\in SelectedParties$ then
17           $\sigma_{id}\leftarrow SigShare_{id}\langle sk_{id},requests\rangle$
18           $reply\langle requests,\sigma_{id}\rangle$
19
20
21
22
23
@@ 25             $\Sigma\leftarrow\{\}$
26
27
28             upon $RandomOrder\langle ID,SelectedParties\rangle$ invocation do
29               $\sigma_{i}$ $\leftarrow$ $CShare\langle r_{id}\rangle$
30               $id=ID.id$
31               $\textbf{send}\langle SHARE,id,\sigma_{id},ID.INSTANCE\rangle$ to all parties
32               wait until $|\Sigma|=2f+1$
33
34              return filterSelectedParties( $CToss\langle r_{id},\Sigma\rangle,SelectedParties$)
35
36
@@ 29               $\sigma_{i}$ $\leftarrow$ $CShare\langle r_{id}\rangle$
39                   $\Sigma\leftarrow{\Sigma\cup\{\sigma_{k}\}}$
40
41
42
Algorithm 2 pVCBC and RandomOrder protocols for party $p_{i}$
4.4 Propose-Recommend
In the pMVBA protocol, the Propose-Recommend step replaces the traditional commit step found in the classical MVBA protocol. This shift is central to enhancing the protocol’s efficiency by significantly reducing complexity. While the classic commit step involves collecting nfn-f verifiable proofs and creating an array of length $n$, our recommend step simplifies the process by broadcasting verifiable proofs directly. The propose step remains largely similar to the one in the classic protocol, but with a key difference: only the selected parties, as determined by the CS process, are responsible for proposing their requests. The input for this step includes a proposal and a threshold-signature, but it does not produce any direct output.
The recommend step then follows, where the input consists of the proposal and the threshold-signature. The output of this step is a list of threshold-signatures (collected from the nfn-f recommendation message), ensuring that at least one verifiable proof reaches the majority of parties. This process is critical for moving the protocol towards an agreement.
The recommend step involves the following key actions:
  • Initialization: A party initializes an empty set $\Sigma$. (line 2)
  • Broadcast Recommendation: A party creates a recommendation message and broadcasts it to all parties. (line 6)
  • Recommendation Collection: Upon receiving a recommendation, a party verifies the threshold-signature and instance. If valid, the party adds the proposal and $\sigma$ to its set $\Sigma$. (lines 11-14)
  • A party waits for nfn-f recommendation messages before proceeding. (line 6-7)
The detailed SCR (Send, Collect, Recommend) protocol is shown in Algorithm 3, which formalizes the steps involved in the recommend process.
1Local variables initialization:
2 $\Sigma\leftarrow\{\}$
3 $CUR\_INSTANCE\leftarrow 1$
4 $rCount\leftarrow 0$
5
6 upon $SCR\langle ID,ID_{k},Proposal,\rho\rangle$ invocation do
7 send $\langle RECOMMENDATION,ID_{k},Proposal,\rho\rangle$ to all parties
8 $CUR\_INSTANCE\leftarrow ID.INSTANCE$
9 $rCount\leftarrow 0$
10 wait until rCount=nfrCount=n-f
11 return $\Sigma$
12
13
14
15upon receiving $\langle RECOMMENDATION,ID_{j},Proposal,\rho\rangle$ from party $p_{j}$ for the first time do
16 if ID.INSTANCE=CUR_INSTANCEID.INSTANCE=CUR\_INSTANCE $\wedge$ $Verify_{id}\langle Proposal,\rho\rangle=true$ $\wedge ID_{j}.id\notin\Sigma$ then
17    $\Sigma\leftarrow{\Sigma\cup\{ID_{j}.id:\langle Proposal,\rho\rangle\}}$
18
19    $rCount\leftarrow rCount+1$
20
21
22
23
Algorithm 3 Recommend Protocol for all parties
4.5 Random Order
Following the Propose-Recommend step, the next crucial phase in the pMVBA protocol is to determine the order in which the selected parties will proceed. The purpose of this Random Order step is to generate a random permutation of the selected parties, ensuring that the process remains fair and secure. This step is particularly important for preventing adversaries from manipulating the message delivery order, which could potentially increase the number of asynchronous rounds required to reach an agreement.
The input for this step includes the party ID and the current protocol instance, and the output is a permutation of the selected parties. The random order is generated after the distribution and receipt of the threshold number of recommendation messages, ensuring the process is not vulnerable to adversarial interference. The Random Order protocol involves the following steps :
  • Coin-Share Generation: Upon invocation of the RandomOrder protocol, a party generates a coin-share $\sigma_{i}$ for the current instance and broadcasts it to every party. The party then waits for $2f+1$ coin-shares. (lines 19-24)
  • Coin-Share Verification: When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share, ensuring it is from $p_{k}$, and accumulates the coin-share in the set $\Sigma$. The party continues to respond to coin-shares until it has received $2f+1$ valid shares. (lines 27-28)
  • Permutation Generation: Once $2f+1$ valid coin-shares have been received, the party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties and filtered the selected parties and return. (line 24 of Algorithm 2).
This process is formalized in the pseudocode for the random-order generation protocol, as shown in Algorithm 2. By generating a random order of the selected parties, this step ensures that the protocol remains robust against adversarial manipulation and maintains the efficiency of the agreement process.
4.6 Sequential-ABBA
The final phase in the pMVBA protocol is the Sequential-ABBA protocol, which is responsible for reaching an agreement on one of the proposals submitted by the selected parties. Building on the random order generated in the previous step, the Sequential-ABBA protocol runs an agreement loop that systematically evaluates each selected party’s proposal. The input to this protocol consists of the permutation list generated in the Random Order step and the list of recommendations obtained from the Propose-Recommend step. The output is the final proposal agreed upon by the parties.
In Sequential-ABBA, we employ an asynchronous binary Byzantine agreement protocol that is biased towards $1$, allowing for efficient agreement on proposals. The protocol iterates through the permutation list, running the agreement process for each selected party until an agreement is achieved on a single proposal. This approach ensures that the final decision is reached in a manner that is both robust and fair, in line with the protocol’s overall design. The detailed construction of the Sequential-ABBA protocol is provided below, where the specific steps and mechanisms are described. This protocol serves as the culminating step in the pMVBA process, ensuring that all parties converge on a single, agreed-upon proposal, thus completing the agreement procedure.
Construction of the Sequential-ABBA
  • The Sequential-ABBA protocol takes two arguments: (i) PermutationList, the permutation of the selected parties, and (ii) RecommendationList, the list of verifiable proofs and proposals the party has received (line 1).
  • Upon invocation of the protocol, a party declares two variables: (i) indexindex, an index number to access the selected parties one by one from the array PermutationList, and (ii) bitbit, initially set to zero to indicate that the parties have not reached any agreement (lines 2-3).
  • While loop: Parties are chosen one after another according to the permutation PermutationList of $\{1,\ldots,f+1\}$. Let indexindex denote the index of the party selected in the current loop (the selected party $p_{index}$ is called the candidate). Each party $p_{i}$ follows these steps for the candidate $p_{index}$ (lines 4-25):
    • Broadcasts a $VOTE$ message to all parties containing $u_{index}=1$ if party $p_{i}$ has received $p_{index}$’s proposal and verifiable proof (including the proposal in the VOTE or from RecommendationList), and $u_{index}=0$ otherwise. (lines 12-18)
    • Waits for $\langle n-f\rangle$ VOTEVOTE messages (line 19) but does not count votes indicating $u_{index}=1$ unless a valid proposal from the party $p_{id}$ has been received—either directly or included in the $VOTE$ message (lines 21 - 22).
    • Runs a binary asynchronous Byzantine agreement biased towards $1$ (see Algorithm 6) to determine whether $p_{index}$ has properly broadcast a valid proposal. Vote $1$ if $p_{i}$ has received a valid proposal from $p_{index}$ through RecommendationList and add the protocol message that completes the verifiable broadcast of $p_{index}$’s proposal to validate this vote. Otherwise, if $p_{i}$ has received nfn-f VOTEVOTE messages containing $u_{index}=0$, then vote 0; no additional information is needed. If the agreement loop decides $1$, exit the loop. (lines 21-25)
  • Upon reaching an agreement, if the proposal is empty, then use the threshold signature $\rho$ to get the proposal from the selected party and return the proposal. (lines 26-28)
  • Upon receiving $\langle ID,VOTE,party,u_{index},m\rangle$, a party checks whether $u_{index^{\prime}}=1$. If it receives $u_{index^{\prime}}=1$, then it assigns $m_{index}=m$. (lines 29-33)
1
[…]
3   $index\leftarrow 0$
4   $bit\leftarrow 0$
5
6  while bit=0 do
7     $\Sigma\leftarrow\{\}$
8
9    $index\leftarrow index+1$
10
11    $proposal\leftarrow\perp$
12     $\rho\leftarrow\perp$
13
14    $id\leftarrow PermutationList[index]$
15     $u_{index}\leftarrow 0$
16     $m_{index}\leftarrow<proposal,\rho>$
@@
18    if $id\in RecommendationList$ then
19
20      $u_{index}\leftarrow 1$
21       $proposal,\rho>\leftarrow RecommendationList[id]$
22       $m_{index}\leftarrow<proposal,\rho>$
23       send $\langle ID,VOTE,party_{id},u_{index},m_{index}\rangle$ to all parties
24
25       else
26         send $\langle ID,VOTE,party_{id},u_{index},\perp\rangle$ to all parties
27
28
29         wait until $|\Sigma|$ =nf=n-f
30
31
32        if $u_{index}=1$ then
33           $v\leftarrow\langle 1,<proposal,\rho>\rangle$
34
35           else
36             $v\leftarrow\langle 0,\perp\rangle$
37
38             $\langle bit,\rho\rangle\leftarrow ABBA\langle v\rangle$ biased towards 11
39
40
41            if $proposal=\perp$ then
42               use $\rho$ to complete the verifiable authenticated broadcast and deliver the proposal. See Appendix A.1
43               return proposal
44
45